For the data set below (a) Determine the least-squares regression line. (b) Compute th x30405060 70 y 113 100 87 85 67 (a) Determine the least-squares regression line (Round to four decimal places as needed)
Answer #1

a) suppose our regression line be ,

y=a+bx ; where b= (X - X(Y-Y (X-X)2

a= \overline{y}-b\overline{x}

by the excel table we have b=(-1070)/(1000) . and a= 90.4-(-1.07 * 50)

= -1.07 =143.9

hence our least squared regression line is ,  y= 143.9+( -1.07)x

b) The sum of the squared residuals for the least squares regression line will be ,

\sum (\hat{y}-a-b*x)^{2}

=46.3

CALCULATIONS >>>>>>>>>>>>>>>>>

X Y X-50 Y-90.4 (X-50)*(Y-90.4) (X-50)^2 Y-143.9+(1.07*X) (Y-143.9+(1.07*X))^2
30 113 -20 22.6 -452 400 1.2 1.44
40 100 -10 9.6 -96 100 -1.1 1.21
50 87 0 -3.4 0 0 -3.4 11.56
60 85 10 -5.4 -54 100 5.3 28.09
70 67 20 -23.4 -468 400 -2 4
x_bar        = 50 y_bar= 90.4 SUM= -1070 SUM= 1000 sum=    46.3